Question 23

25.00 cm3 of a 0.7892 mol.dm-3 solution of potassium hydroxide is transferred to an empty 350.00 cm3 volumetric flask. This flask is made up to the mark with distilled water and then shaken well. The concentration of the potassium hydroxide in this second flask is:

a. 0.01109 mol.dm-3

b. 0.05637 mol.dm-3

c. 0.7892 mol.dm-3

d. 17.74 mol.dm-3

e. None of the above.

Question 24

What is the molarity of glacial acetic acid (CH3COOH, Mr = 60.05 g/mol) at 25°C given that the density of acetic acid at that temperature is 1.049 g/mL?

a. 0.0174 M

b. 0.057 M

c. 57.2 M

d. 60.1 M

e. None of the above

 Question 25

Calculate the molarity of a KF (aq) solution containing 116.2 g of KF in 3.00 L of solution?

a. 0.02585 M

b. 0.667 M

c. 38.7 M

d. 58.1 M

e. None of the above


1
Expert's answer
2022-06-07T10:50:03-0400

23.

n(KOH)= CM * V = 0,7892 * 0,025 = =0,01973mol

CM = n/V = 0,01973mol / 0,35 l= 0,05637 M

Answer: B

24.

V=m/p=60,05/1,049 =57,245 ml =0,057245 l

CM = n / V = 1mol / 0,057245 l = 17,47 mol/l

Amswer: E

25.

M(KF) = 58 g/mol

n = m / M = 116,2 / 58 = 2mol

CM = n / V = 2mol / 3 l = 0,667 M

Answer: B